OpenStudy (nuccioreggie):

Simplify negative 2 over 3 ÷ negative 7 over 4. negative 7 over 6 negative 8 over 21 8 over 21 7 over 6

OpenStudy (nuccioreggie):

@mathmath333

OpenStudy (mathmath333):

\(\large \color{black}{\begin{aligned} \dfrac{-2}{3} \div \dfrac{-7}{4}\hspace{.33em}\\~\\ \end{aligned}}\)

OpenStudy (mathmath333):

\(\large \color{black}{\begin{aligned} \dfrac{-2}{3} \div \dfrac{-7}{4}\hspace{.33em}\\~\\ =\dfrac{-2}{3} \times \dfrac{4}{-7}\hspace{.33em}\\~\\ \end{aligned}}\)